# Enhancing Action Recognition by Leveraging the Hierarchical Structure of Actions and Textual Context

### v1

@@ Abstract
The sequential execution of actions and their hierarchical structure consisting of different levels of abstraction, provide features that remain unexplored in the task of action recognition. In this study, we present a novel approach to improve action recognition by exploiting the hierarchical organization of actions and by incorporating contextualized textual information, including location and prior actions to reflect the sequential context. To achieve this goal, we introduce a novel transformer architecture tailored for action recognition that utilizes both visual and textual features. Visual features are obtained from RGB and optical flow data, while text embeddings represent contextual information. Furthermore, we define a joint loss function to simultaneously train the model for both coarse and fine-grained action recognition, thereby exploiting the hierarchical nature of actions. To demonstrate the effectiveness of our method, we extend the Toyota Smarthome Untrimmed (TSU) dataset to introduce action hierarchies, introducing the Hierarchical TSU dataset. We also conduct an ablation study to assess the impact of different methods for integrating contextual and hierarchical data on action recognition performance. Results show that the proposed approach outperforms pre-trained SOTA methods when trained with the same hyperparameters. Moreover, they also show a 17.12% improvement in top-1 accuracy over the equivalent fine-grained RGB version when using ground-truth contextual information, and a 5.33% improvement when contextual information is obtained from actual predictions.
keywords:
Action Recognition , Action Hierarchies , Action Context , Vision-Language Transformer
journal: Image and Vision Computing
[…]
[label1]organization=Department of Computer Technology, University of Alicante, city=Alicante, country=Spain
[…]
[label2]organization=Institute of Computer Science, FORTH, city=Heraklion, state=Crete, country=Greece
1 Introduction
Action recognition has become a fundamental task in a wide range of real-world applications. It involves the understanding of actions occurring within video sequences, allowing for diverse applications in everyday scenarios such as video surveillance [1], autonomous driving [2, 3], and human-robot interaction [4].
@@ 1 Introduction
In some of these cases, the effectiveness of the provided solutions depends on their ability to operate online and in real-time. However, in several other cases (e.g. routine analysis [5, 6]), offline analysis of behavior is sufficient. In such cases, the consideration of the broader context provided by the entire video, may lead to improved model performance.
@@
In the offline setting, action recognition and action detection emerge as the primary tasks in analyzing trimmed and untrimmed videos, respectively. Action recognition involves categorizing videos into predefined classes [7, 8], whereas action detection extends this by identifying the temporal boundaries of each action within the video [9, 10]. While the former pursues a straightforward goal, the latter benefits from the sequential nature of actions to enrich the temporal context, thereby reinforcing the robustness of the model.
@@
The definition of the different action categories is typically left to the discretion of dataset authors, resulting in varying levels of abstraction. For example, the task of making coffee may be decomposed into several steps involving different objects: open coffee maker, place filter, etc. While the coarse granularity is sufficient in some cases, certain applications may require additional refinement to achieve specific goals. For example, robotic applications often need more detailed annotations, including per-joint instructions.
@@
This work presents an action recognition architecture that exploits both the hierarchical structure and the sequential execution of actions. To this end, we adopt a transformer encoder built upon state-of-the-art methods [11, 12, 13] that use this architecture to model temporal dependencies. By exploiting the self-attention mechanism inherent in transformer architectures, this model exceeds previous methods such as 3D CNNs [14, 15, 16] and RNNs [17, 18, 19], which often face challenges in capturing long-range temporal dependencies.
@@
The Transformer architecture [20], originally designed for natural language processing, has undergone numerous improvements in recent years [21, 22, 23]. These improvements have enabled the models to handle temporal dependencies more effectively, and capture contextual information within textual data. We explore the benefits of these models to incorporate contextual data for action recognition in two different ways: (1) evaluating the effect of rephrasing to enrich textual descriptions of past actions along with camera location information leveraging instructional models such as GPT-3.5 [21] and Llama 3 [24]; and (2) encoding textual data into feature vectors using transformer text encoders [23, 25]. By fusing contextual information in the form of text embeddings with visual features prior to classification, we improve the overall representation of the ongoing action, enabling for more accurate action recognition.
@@
Finally, we train our model with a dual objective: coarse-grained and fine-grained action classification. The rationale behind this approach is rooted in the notion that learning together the coarse- and the fine-grained action facilitates the recognition of both. Previous studies [26, 27] have explored action hierarchies under the premise that hyperbolic geometry naturally encodes hierarchical structures. Other works [28, 29] exploit the goal of the video, i.e. a coarse-grained action, to improve action anticipation performance. In our method, we propose a joint loss function to allow the model to better discriminate fine-grained actions. For experimentation, we extend the existing Toyota Smarthome Untrimmed (TSU) dataset [30], incorporating action hierarchies into Activities of Daily Living (ADL). Code and extended annotations resulting in the Hierarchical TSU dataset will become publicly available on Github [^1] https://github.com/3dperceptionlab/HierarchicalActionRecognition.
@@
  • 1.
    We introduce a novel vision-language transformer architecture that exploits hierarchical and contextual information from actions to improve accuracy in the task of action recognition. Our approach outperforms state-of-the-art visual-only models.
  • 2.
    We present the Hierarchical TSU dataset, an extension of the original TSU dataset [30], incorporating coarse-grained annotations and enriched textual information. Additionally, we compare various strategies for generating textual descriptions of contextual data by leveraging previous actions and camera location.
  • 3.
    We conduct extensive experiments to demonstrate the effectiveness of our approach and to compare it with relevant state of the art methods. Additionally, we present detailed ablation studies that analyze the contributions of our model’s components, as well as the impact of hierarchical and contextual information.
The remainder of the paper is structured as follows. Section 2 summarizes the latest related works. In Section 3 we describe in detail the proposed method. The experiments and results are reported in Section 4, followed by an ablation study in Section 5. Finally, the main conclusions of this work are presented in Section 6.
2 Related Work
We present an overview of pertinent literature concerning action recognition, focusing on the integration of language models and hierarchical structures of actions as tools to enhance performance.
2.1 Action Recognition
@@
Second, temporal modeling transformers that use pre-trained feature extractors from large datasets such as Kinetics-400 [14] and ImageNet [39]. OadTR [11] focus on temporal modeling using decoded RGB frames and frozen frame-level feature extractors. In addition, optical flow is computed from the RGB data to improve accuracy. Similarly, TIM [40] uses frozen visual and audio encoders for feature extraction. Features include a timestamp provided by a Time Interval MLP, so that the model can be queried about the events at a given interval in a specific modality. Other approaches incorporate the vision transformer, ViT [41], to fine-tune the feature extractor on the corresponding datasets [7, 8, 42], yielding better results with a remarkable increase in terms of computing cost.
@@
2.2 Language Models for Action Recognition
The remarkable capabilities of large language models in temporal modeling and feature representation have been thoroughly investigated [21, 22, 23]. These capabilities have found application in various vision tasks, notably in image and video captioning. While these approaches focus on generating textual descriptions of videos [48], recent research has explored the use of language models to improve action recognition results. For example, ActionCLIP [7] uses a contrastive learning approach inspired by CLIP [49] to improve action recognition. Instead of image captions, action classes are converted to prompts, which are then compared to the aggregated representation of a video. Spatial features from frames are extracted from a fine-tuned version of CLIP’s visual encoder, based on ViT [41]. Similarly, VideoCLIP [50] extends this learning paradigm to various video understanding tasks. Building on these advances, Text4Vis [8, 42] proposes to initialize a frozen classifier for action recognition using text embeddings derived from language models. Moreover, BIKE [51] introduces a novel framework that facilitates bi-directional cross-model knowledge transfer from vision to language models, with the aim of improving action recognition in videos.
Furthermore, the capabilities of language models in capturing temporal dependencies and contextual understanding have been exploited to model past actions as well [52]. VLMAH [53] presents a visual-linguistic approach to modeling action history that is particularly useful for instructional videos due to the sequential nature of the actions. Similarly, Furnari and Farinella [54] use “rolling-unrolling” LSTMs to succinctly summarize past actions. In contrast, AntGPT [29] exploits large language models for in-context learning [21] long-term action anticipation, i.e. by providing few ground-truth examples. Similarly, [55] explores the benefits of incorporating object representations for this task. Encoded representations of cropped objects from RGB frames, along with bounding boxes and object labels encoded as text embeddings, are shown to improve the model’s performance.
2.3 Action Hierarchies
@@
In addition, Timeception [57] redefines the notion of activity, restricting it to “complex actions” characterized by: (1) composition - consisting of several simpler actions, (2) temporal order of these actions, and (3) extent - recognizing the variability in temporal length between actions. By introducing the Timeception layer, the architecture tolerates both long-range temporal dependencies and variations in the temporal extent.
@@
The previously introduced AntGPT [29] also benefits from hierarchical information by incorporating goal information extracted from past actions using a large language model. The resulting text embedding is fused with visual observations to improve long term action anticipation. Similarly, in [28], authors exploit the goal concept to improve action anticipation in industrial scenarios by introducing a consistency loss to ensure alignment between coarse-grained and fine-grained predictions.
@@
It is important to note the scarcity of annotated hierarchical data in the ADL domain, which prompts researchers in [26] to manually extend existing datasets for experimentation. Alternatively, synthetic data generators, such as Virtualhome [58], offer a hierarchical approach to generating video data, providing a valuable resource to address this limitation.
3 Methodology
We present the proposed method for action recognition using hierarchical action structures and textual context. Figure 1 illustrates an overview of the proposed architecture.
@@ 3 Methodology
Figure 1: Overview of the proposed action recognition architecture. From a video input, frozen feature extractors obtain spatial features from RGB frames and temporal motion features from optical flow. These features, along with a class token, are processed through individual video transformer encoders for each visual modality, capturing long-range temporal dependencies. Dashed lines indicate the use of either RGB or flow features and their respective embeddings, though only a single video encoder is depicted for simplicity. Additionally, DistilBERT extracts textual features that represent the current location and previous actions. Coarse-grained actions are recognized based on contextual information, specifically utilizing the textual embeddings while the fine-grained classifier leverages the fused features from the fusion transformer.
3.1 Video Encoder
@@
We adopt a standard Transformer’s encoder to model temporal dependencies as in [11]. Given a trimmed video $V=\{f_{t}\}^{0}_{t=-T}$ associated with a coarse-grained action, and one or more fine-grained actions, the frames $f_{t}$ are grouped into blocks of size $B$, resulting in $(T+1)/B$ chunks. For each block, we use a frozen feature extractor to obtain spatial features from RGB and optical flow frames, and employ separate video encoders with the same structure for each modality. We use the central frame of the block for RGB features, and the $B$ frames sequence for optical flow features. These features are mapped into a $D$ dimensional feature space, which we can formalize as $F=\{\mathcal{O}_{t}\}^{0}_{t=-T}\in\mathbb{R}^{(T+1)\times D}$, where $\mathcal{O}_{t}$ represents the output token of the transformer encoder at time $t$.
@@
where $Q_{i}=X^{\prime}W_{i}^{q}$, $K_{i}=X^{\prime}W_{i}^{k}$ and $V_{i}=X^{\prime}W_{i}^{v}$ are linear layers applied to the input sequence, $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{D\times\frac{D}{N_{head}}}$ with $N_{head}$ represent the number of heads, and $\frac{1}{\sqrt{d_{k}}}$ is a scaling factor that helps stabilizing the training and speed up convergence. Note that queries, keys and values are all vector representations. The outputs of each head are concatenated and fed into a linear layer as:
@@
where $n\in N$ is the n-th encoder layer with a total of $N$, $m^{\prime}$ denotes the output of the first element-wise addition in the encoder layer and $m_{N}\in\mathbb{R}^{(T+2)\times D}$ is the final feature representation of the last encoder layer. For the remaining of the paper, and for simplicity, we use $m_{CLS}\in\mathbb{R}^{D}$ to denote the output representation of the class token from an encoder. This notation is extended to $m_{RGB}$ and $m_{Flow}$ to refer to the same token of the corresponding modality.
3.2 Contextualized Textual Information
Large language models have shown exceptional abilities in temporal modeling and understanding context. To leverage these capabilities, we propose an approach that models past information to enhance the contextualization of current action recognition. Specifically, given the location and $N$ previous actions, we use a prompt (as illustrated in Figure 1) to provide context about prior events, thus modeling longer temporal dependencies. From the generated sentences, we extract feature vectors using DistilBERT [25], a distilled variant of BERT [23], which employs deep bidirectional transformer encoders for text comprehension. As for the video encoder, we utilize the class token to represent the entire input sequence, denoted as $m_{Txt}$.
In addition to DistilBERT, we compare its performance against BERT and explore the use of LLMs for rephrasing to increase variability and enrich the data. Section 5 presents results for different setups, including no rephrasing, and rephrasing using GPT-3.5 [21] and Llama 3 [24], as well as an analysis of the optimal number of past actions and the impact of location information on contextualization.
3.2.1 Exploiting Hierarchical Information from Context

### v2

@@ Abstract
We propose a novel approach to improve action recognition by exploiting the hierarchical organization of actions and by incorporating contextualized textual information, including location and previous actions, to reflect the action’s temporal context. To achieve this, we introduce a transformer architecture tailored for action recognition that employs both visual and textual features. Visual features are obtained from RGB and optical flow data, while text embeddings represent contextual information. Furthermore, we define a joint loss function to simultaneously train the model for both coarse- and fine-grained action recognition, effectively exploiting the hierarchical nature of actions. To demonstrate the effectiveness of our method, we extend the Toyota Smarthome Untrimmed (TSU) dataset by incorporating action hierarchies, resulting in the Hierarchical TSU dataset, a hierarchical dataset designed for monitoring activities of the elderly in home environments. An ablation study assesses the performance impact of different strategies for integrating contextual and hierarchical data. Experimental results demonstrate that the proposed method consistently outperforms SOTA methods on the Hierarchical TSU dataset, Assembly101 and IkeaASM, achieving over a 17% improvement in top-1 accuracy.
I Introduction
Recognizing actions in video sequences has become a fundamental task in a wide range of real-world applications, particularly in assistive technologies that enhance safety, efficiency, and quality of life. These technologies are increasingly integral to diverse domains, including video surveillance [1], autonomous driving [2, 3], and human-robot interaction [4]. Specifically, video monitoring systems, such as smarthomes and robots, stand to benefit immensely from these advances. These systems have the potential to significantly improve the quality of life for older adults in domestic environments and enhance worker safety in industrial settings, thereby driving innovation in both healthcare and industrial safety.
@@ I Introduction
In some of these cases, the effectiveness of the provided solutions depends on their ability to operate online and in real-time. In other scenarios, however, offline analysis of behavior may be sufficient. In such cases, considering the broader context provided by the entire video can improve model performance. For example, routine analysis can be used to detect behavioral changes in elderly individuals [5, 6], or to identify errors in industrial assembly tasks [7, 8, 9].
@@
In the offline setting, action recognition and action detection emerge as the primary tasks in analyzing trimmed and untrimmed videos, respectively. Action recognition involves categorizing videos into predefined classes [10, 11], whereas action detection extends this by identifying the temporal boundaries of each action within the video [12, 13]. While the former pursues a straightforward goal, the latter benefits from the sequential nature of actions to enrich the temporal context, thereby reinforcing the robustness of the employed model and solution.
@@
A key aspect of both tasks is the granularity of the action categories, which can significantly affect the performance and applicability of the model. The definition of the different action categories is typically left to the discretion of the dataset authors, resulting in varying levels of abstraction. For example, the task of make coffee may be decomposed into several steps involving different objects: open coffee maker, place filter, etc. While the coarse granularity is sufficient in some cases, certain applications may require additional refinement to achieve specific goals. For example, robotics applications often require more detailed annotations, including per-joint instructions.
@@
This work presents an action recognition architecture that exploits both the hierarchical structure and the sequential execution of actions. To this end, we adopt a transformer encoder built on the basis of state-of-the-art methods [14, 15, 16] that use this architecture to model temporal dependencies. By exploiting the self-attention mechanism inherent in transformer architectures, this model exceeds previous methods such as 3D CNNs [17, 18, 19] and RNNs [20, 21, 22], which often face challenges in capturing long-range temporal dependencies.
@@
The Transformer architecture [23], originally designed for natural language processing, has undergone numerous improvements in recent years [24, 25, 26]. These improvements have enabled the models to handle temporal dependencies more effectively and to capture contextual information within textual data. We explore the benefits of these models to incorporate contextual data for action recognition in two different ways. First, by evaluating the effect of rephrasing to enrich textual descriptions of past actions along with camera location information which is achieved by leveraging instructional models such as GPT-3.5 [24] and Llama 3 [27]. Second, by encoding textual data into feature vectors using transformer text encoders [26, 28]. By fusing contextual information in the form of text embeddings with visual features prior to classification, we improve the overall representation of the ongoing action, enabling more accurate action recognition.
@@
Our model is trained with the dual objective of both coarse-grained and fine-grained action classification. The rationale behind this approach is rooted in the idea that learning both coarse-grained and fine-grained actions together facilitates the recognition of both. Previous studies [29, 30] have explored action hierarchies under the premise that hyperbolic geometry naturally encodes hierarchical structures. Other works [31, 32] exploit the goal of the video, i.e. a coarse-grained action, to improve action anticipation performance. In our method, we propose a joint loss function to allow the model to better discriminate fine-grained actions.
For experimentation, we extend the existing Toyota Smarthome Untrimmed (TSU) dataset [33], incorporating action hierarchies into Activities of Daily Living (ADL). The resulting dataset, the Hierarchical TSU dataset, is, to the best of our knowledge, the first hierarchical dataset for ADL. Additionally, we evaluate the proposed method on the Assembly101 [8] and IkeaASM [34] action recognition benchmarks to provide further insights on the benefits of incorporating contextual information and action hierarchies.
Our implementation and the extended annotations for the Hierarchical TSU dataset are publicly available on Github [^1] https://github.com/3dperceptionlab/HierarchicalActionRecognition.
@@
  • We present the Hierarchical TSU dataset, an extension of the original TSU dataset [33], that incorporates coarse-grained annotations and contextual information. Additionally, we compare various strategies for generating textual descriptions of contextual data by leveraging previous actions and camera location.
  • We introduce a novel vision-language transformer architecture for action recognition that leverages hierarchical and contextual information, leading to more accurate action classification. Our approach outperforms state-of-the-art action recognition models in Assembly101 and IkeaASM benchmarks and serves as a strong baseline for the Hierarchical TSU dataset.
  • We conduct extensive experiments to demonstrate the effectiveness of our approach. In addition to outperforming existing methods, we present detailed ablation studies that analyze the contributions of each component, as well as the impact of hierarchical and contextual information, including the role of large language models and the influence of the number of past actions, and of location data.
The remainder of the paper is structured as follows. Section II summarizes the latest related work. In Section IV we describe in detail the proposed method. The experiments and results, including details on the Hierarchical TSU dataset, are reported in Section V, followed by the ablation study in Section VI. Finally, the main conclusions of this work are discussed in Section VII.
II Related Work
We present an overview of pertinent literature concerning action recognition, focusing on the integration of language models and hierarchical structures of actions as tools to enhance action recognition performance.
II-A Action Recognition
@@
The second category comprises temporal modeling transformers that use pre-trained feature extractors from large datasets such as Kinetics-400 [17] and ImageNet [43]. OadTR [14] focus on temporal modeling using decoded RGB frames and frozen frame-level feature extractors. In addition, optical flow is computed on the RGB data to improve accuracy. Similarly, TIM [44] uses frozen visual and audio encoders for feature extraction. Features include a timestamp provided by a Time Interval MLP, so that the model can be queried about the events at a given interval in a specific modality. Other approaches incorporate the vision transformer, ViT [45], to fine-tune the feature extractor on the corresponding datasets [10, 11, 46], yielding better results with a remarkable increase in terms of computing cost.
@@
II-B Language Models for Action Recognition
The remarkable capabilities of large language models in temporal modeling and feature representation have been thoroughly investigated in recent years [24, 25, 26]. These capabilities have found application in various vision tasks, notably in image and video captioning. These approaches focus on generating textual descriptions of videos [52]. Recenttly, research has explored the use of language models to improve action recognition results. For example, ActionCLIP [10] uses a contrastive learning approach inspired by CLIP [53]. Instead of image captions, action classes are converted to prompts, which are then compared to the aggregated representation of a video. Spatial features from frames are extracted from a fine-tuned version of CLIP’s visual encoder, based on ViT [45]. Similarly, VideoCLIP [54] extends this learning paradigm to various video understanding tasks. Building on these advances, Text4Vis [11, 46] proposes to initialize a frozen classifier for action recognition using text embeddings derived from language models. Moreover, BIKE [55] introduces a novel framework that facilitates bi-directional cross-model knowledge transfer from vision to language models, with the aim of improving action recognition in videos.
Furthermore, the capabilities of language models in capturing temporal dependencies and contextual understanding have been exploited to model past actions as well [56]. VLMAH [57] presents a visual-linguistic approach to modeling action history that is particularly useful for instructional videos due to the sequential nature of the actions. Similarly, Furnari and Farinella [58] use “rolling-unrolling” LSTMs to succinctly summarize past actions. In contrast, AntGPT [32] exploits large language models for in-context learning [24] long-term action anticipation, i.e. by providing few ground-truth examples. Similarly, Zhang et al. [59] explore the benefits of incorporating object representations for this task. Encoded representations of cropped objects from RGB frames, along with bounding boxes and object labels encoded as text embeddings, are shown to improve the model’s performance.
II-C Action Hierarchies
Hierarchical structures have been explored in various ways for action understanding tasks, such as modeling features at multiple levels [13, 60, 61, 62] or leveraging annotations with varying granularity. In this work, we focus on the latter to investigate the potential of structured datasets for improving action understanding.
@@ II-C Action Hierarchies
Alternatively, Timeception [64] redefines the notion of activity, restricting it to “complex actions” characterized by: (1) composition - consisting of several simpler actions, (2) temporal order of these actions, and (3) extent - recognizing the variability in temporal length between actions. By introducing the Timeception layer, the architecture tolerates both long-range temporal dependencies and variations in the temporal extent.
@@
The previously introduced AntGPT [32] also benefits from hierarchical information by incorporating goal information extracted from past actions using a large language model. A large language model is used to extract these goals in the form of text embeddings, which are then fused with visual observations to enhance long-term action anticipation. To generate hierarchical labels, AntGPT relies on twelve in-context examples, either manually curated or pseudo-labeled using large language models applied to video titles or descriptions. Similarly, in [31], authors exploit the goal concept to improve action anticipation in industrial scenarios by introducing a consistency loss to ensure alignment between coarse-grained and fine-grained predictions.
The industrial domain has particularly benefited from hierarchical annotations [8, 34], which may be explicitly provided or derived by defining coarse-grained related goals (e.g., assembling or disassembling a product). Other domains, such as autonomous driving, have also explored hierarchical annotations to capture the progression from high-level driving maneuvers to specific low-level actions [65].
@@
Despite these advances, it is important to note the scarcity of annotated hierarchical data in the ADL domain, which prompts researchers in [29] to manually extend existing datasets for experimentation. Alternatively, synthetic data generators, such as Virtualhome [66], offer a hierarchical approach to generating video data, providing a valuable resource to address this limitation.
III Hierarchical TSU Dataset
Figure 1: Hierarchical annotation example in the TSU dataset. Video frames are annotated with fine-grained action labels, including composite activities (e.g., Make coffee). Each color corresponds to one of the coarse-grained action categories.
We introduce the Hierarchical TSU dataset, an extension of the Toyota Smarthome Untrimmed dataset [33], designed for the study of action hierarchies in activities of daily living. This dataset aims to incorporate and evaluate two complementary sources of information for action recognition: (1) contextual cues, such as prior actions, spatial location, and scene context, and (2) hierarchical organization of actions, reflecting the natural structure of human behavior.
We adopt the TSU dataset as the foundation for our work due to several key advantages. Unlike other ADL datasets (e.g., Toyota Smarthome [67] or ETRI-Activity [68]), TSU provides temporally ordered annotations of actions, static camera viewpoints with known spatial locations, and a comprehensive taxonomy of fine-grained action labels. Specifically, it includes 51 distinct action classes recorded from 7 fixed cameras in various rooms of a house. The dataset features 536 untrimmed videos, each averaging 21 minutes in duration, and captures daily activities performed by 18 elderly individuals aged 60 to 80. This diversity in subject demographics and household settings makes TSU particularly well-suited for hierarchical and context-aware action recognition. In total, TSU contains over 64,000 annotated fine-grained action segments, which correspond to the trimmed action segments that compose the Hierarchical TSU dataset.
Among the 51 annotated actions are elementary activities such as walk and get water, as well as composite actions like make coffee and cook, which often encompass or temporally overlap with elementary actions. Although these composite and elementary actions exhibit clear hierarchical relationships, we preserve the original flat annotation structure of the TSU dataset to maintain consistency. Building upon this, we introduce an additional semantic layer by organizing the 51 action classes into a two-level hierarchy. Specifically, we manually assign each fine-grained action to one of seven coarse-grained categories, based on contextual semantics, location of execution, and visual similarity. The resulting high-level categories are: beverage preparation, general household activities, cleaning, prepare breakfast, use household appliances, cook, and drink. Table I summarizes the distribution of fine-grained actions within each coarse-grained category, as well as the number of annotated segments per class. While our work focuses on action recognition, we note that the introduced hierarchical annotations are also well-suited for action detection tasks within the TSU setup, as they enable richer temporal modeling and more structured predictions.
TABLE I: Distribution of fine-grained actions per coarse-grained category.
Coarse-Grained # Fine-Grained Samples Avg. Dur. (s)
Beverage Preparation 11 2263 6.4
General Household Activities 14 40779 7.8
Cleaning 5 3457 10.1
Prepare breakfast 5 3787 7.83
Use Household Appliances 7 5645 13.33
Cook 5 6242 7.83
Drink 4 5241 2.87
For contextual action recognition, we extract trimmed segments corresponding to individual fine-grained actions from the untrimmed videos. Because composite activities may span or overlap multiple fine-grained segments, the dataset includes overlapping annotations at both hierarchical levels, as illustrated in Figure 1. To enrich the temporal context, we additionally process preceding actions in the sequence, encoding them as textual input. Finally, the known spatial configuration of the cameras, detailed in [33], is leveraged to associate each action with its location, thereby providing valuable spatial context for downstream models.
IV Methodology
@@
Figure 2: Overview of the proposed action recognition architecture. From a video input, frozen feature extractors obtain spatial features from RGB frames and temporal motion features from optical flow. These features, along with a class token, are processed through individual video transformer encoders for each visual modality, capturing long-range temporal dependencies. Dashed lines indicate the use of either RGB or flow features and their respective embeddings, though only a single video encoder is depicted for simplicity. Additionally, DistilBERT extracts textual features that represent the current location, previous actions and dataset context. Coarse-grained actions are recognized based on contextual information, specifically utilizing the textual embeddings while the fine-grained classifier leverages the fused features from the fusion transformer.
We present the proposed method for action recognition using hierarchical action structures and contextual information. Figure 2 illustrates an overview of the proposed architecture.
IV-A Video Encoder
We adopt a standard Transformer’s encoder to model temporal dependencies as in [14]. As noted earlier, the self-attention mechanism in transformer architectures has significantly outperformed previous methods based on 3D CNNs and RNNs, especially in the context of temporal modeling for videos. This mechanism allows for more effective capture of long-range dependencies and complex temporal relationships within video sequences.
@@ IV-A Video Encoder
Given a trimmed video $V=\{f_{t}\}^{0}_{t=-T}$ associated with a coarse-grained action, and one or more fine-grained actions, the frames $f_{t}$ are grouped into blocks of size $B$, resulting in $(T+1)/B$ chunks. For each block, we use a frozen feature extractor to obtain spatial features from RGB and optical flow frames, and employ separate video encoders with the same structure for each modality. We use the central frame of the block for RGB features, and the $B$ frames sequence for optical flow features. These features are mapped into a $D$ dimensional feature space, which we can formalize as $F=\{\mathcal{O}_{t}\}^{0}_{t=-T}\in\mathbb{R}^{(T+1)\times D}$, where $\mathcal{O}_{t}$ represents the output token of the transformer encoder at time $t$.
@@
where $Q_{i}=X^{\prime}W_{i}^{q}$, $K_{i}=X^{\prime}W_{i}^{k}$ and $V_{i}=X^{\prime}W_{i}^{v}$ are linear layers applied to the input sequence, $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{D\times\frac{D}{N_{head}}}$ with $N_{head}$ represent the number of heads, and ${1}/{\sqrt{d_{k}}}$ is a scaling factor that helps stabilizing the training and speed up convergence. Note that queries, keys and values are all vector representations. The outputs of each head are concatenated and fed into a linear layer as:
@@
where $n\in N$ is the n-th encoder layer with a total of $N$, $m^{\prime}$ denotes the output of the first element-wise addition in the encoder layer, and $m_{N}\in\mathbb{R}^{(T+2)\times D}$ is the final feature representation of the last encoder layer. For the remaining of the paper, and for simplicity, we use $m_{CLS}\in\mathbb{R}^{D}$ to denote the output representation of the class token from an encoder. This notation is extended to $m_{RGB}$ and $m_{Flow}$ to refer to the same token of the respective modality.
IV-B Contextualized Textual Information
Large language models have shown exceptional abilities in temporal modeling and context understanding. To leverage these capabilities, we propose an approach that models past information to enrich the contextualization of ongoing actions, thereby improving action recognition performance. Specifically, given the location and $N$ previous actions, we use a prompt (as illustrated in Figure 2) to provide contextual information, including dataset domain (e.g. an elderly person), environment (e.g. camera location) and prior events, thus modeling longer temporal dependencies. From the generated sentences, we extract feature vectors using DistilBERT [28], a distilled variant of BERT [26], which employs deep bidirectional transformer encoders for text comprehension. As for the video encoder, we utilize the class token to represent the entire input sequence, denoted as $m_{Txt}$.
In addition to DistilBERT, we compare its performance against BERT and explore the use of LLMs for rephrasing to enhance data variability and richness. Section VI presents results for different setups, including no rephrasing, and rephrasing using GPT-3.5 [24] and Llama 3 [27], as well as an analysis of the optimal number of past actions and the influence of location information on contextualization.
IV-B1 Exploiting Hierarchical Information from Context